a company estimates that it has a 30% chance of being successful in bidding on a $40,000 contract. If it costs $6,000 in consultant fees to prepare the bid,what is the expected gain or loss if it decides to bid on this contract

Answer by jim_thompson5910(35256)

Let,
A = probability of getting contract
B = net proceeds if contract is obtained
C = probability of not getting contract
D = net proceeds if contract is not obtained


A = 0.30
B = 40000 - 6000 = 34000
C = 1-A = 1-0.30 = 0.70
D = 0-6000 = -6000


expected value = A*B + C*D
expected value = 0.3*34000 + 0.7*(-6000)
expected value = 6000

The expected value is $6,000
This means the expected gain is $6,000
It is a gain because of the positive expected value.